OpenStudy (defiance):

Will medal... 2x-9/x^2- x- 6 = A/x-3 + B/x+2 / as in fraction and ^2 as squared

OpenStudy (misty1212):

HI!!

OpenStudy (misty1212):

partial fractions?

OpenStudy (defiance):

Mhm.

OpenStudy (defiance):

@ganeshie8

OpenStudy (defiance):

Hello.?

OpenStudy (defiance):

Ganeshie8, can you help?

ganeshie8 (ganeshie8):

you may use the cover up method http://www.math-cs.gordon.edu/courses/ma225/handouts/heavyside.pdf

ganeshie8 (ganeshie8):

\[\dfrac{2x-9}{x^2- x- 6} = \dfrac{A}{x-3} + \dfrac{B}{x+2} \] \[\dfrac{2x-9}{(x-3)(x+2)} = \dfrac{A}{x-3} + \dfrac{B}{x+2} \]

ganeshie8 (ganeshie8):

Look at the "left hand side" expression. Cover up \(x-3\) and plugin \(x=3\), what do you get ?

ganeshie8 (ganeshie8):

\[\dfrac{2x-9}{(x+2)} \] plugin x=3 above

OpenStudy (defiance):

2(3)-9/5

OpenStudy (defiance):

So, 1 / 5?

ganeshie8 (ganeshie8):

(2(3)-9)/5 (6-9)/5 = ?

OpenStudy (defiance):

1/5 and -3/5

ganeshie8 (ganeshie8):

what 1/5 ?

OpenStudy (defiance):

Wait sorry.

OpenStudy (defiance):

i get -3/5 on both

ganeshie8 (ganeshie8):

Lets work it step by step, shall we ? so far we have \(A=-\dfrac{3}{5}\)

OpenStudy (defiance):

Yes.

ganeshie8 (ganeshie8):

Next, cover up \(x+2\) and plugin \(x=-2\) in : \[\dfrac{2x-9}{(x-3)} \]

OpenStudy (defiance):

13/5?

ganeshie8 (ganeshie8):

Yep! you may double check with wolfram http://www.wolframalpha.com/input/?i=partial+fractions+%282x-9%29%2F%28x%5E2-+x-+6%29

OpenStudy (defiance):

Thanks so much!

ganeshie8 (ganeshie8):

yw